A compressed air vessels has inside diameter of 2m and a plate of 12.5mm.if tensile stress induced in plating is to be limited to 100mpa,calculate for both the longitudinal and circumferential seams the pre




1
Expert's answer
2021-04-01T01:11:09-0400

Diameter of vessel is 2m, and thickness of plate is 12.5 mm "= 12.5 \times 10 ^{-3} m"

Longitudinal stress "= \sigma _l= \frac{pd}{2t \eta_c}=\frac{100\times 2}{4\times 12.5\times 10^{-3}}=" 4000Mpa

and circumferential stress

"\sigma _c= \frac{pd}{2t \eta_l}=\frac{100\times 2}{2\times 12.5\times 10^{-3}}=" 8000 Mpa